no load voltage of power source is 220 volt.it provides power of 20 watt in a load.if current value is 0.4 ampere then maximum power supply of power source
1
Expert's answer
2020-08-02T15:11:12-0400

Maximum power transfer theorem states that maximum power is delivered to a load when the load resistance is equal to the source resistance.


Therefore we calculate the load resistance using the formula;


"P=I^{2}R"


"20=0.4^{2}R"


R= 50 ohms


using the maximum power transfer formula


"\frac{V^2}{4\times R}" ,


maximum power supplied by the source is


"\frac{220}{4\times50}=242"


Answer =242